[figure: 9.7 cm]

MAGNETIC PROPERTIES OF THE RRhGe (R=Gd-Tm) COMPOUNDS

W. Bażela^{1}, M. Hofmann^{2}, B. Penc^{3}, A. Szytuła^{3} and A. Zygmunt^{4} 1 Institute of Physics, Technical University, 30-084 Kraków, Pochor_{ι} ażych 1, Poland 2 Berlin Neutron Scattering Centre, Hahn-Meitner Institute, Berlin-Wannsee, Germany 3 Institute of Physics, Jagiellonian University, Reymonta 4, 30-059 Kraków, Poland

4 W. Trzebiatowski Institute of Low Temperatures and Structural Research, Polish Academy of Sciences, Wrocław, Poland

The magnetic properties of the ternary intermetallics RRhGe (R=Gd-Tm) have been investigated by magnetic and neutron diffraction measurements. These compounds crystallize in a TiNiSi-type orthorhombic structure and are antiferromagnets with the Néel temperature 30.8 K for R=Gd, 23 K (Tb), 20 K (Dy), 4.6 K (Ho), 9.5 K (Er) and 6.2 K (Tm). The neutron diffraction data reveal that these compounds exhibit different magnetic structures: sine modulated for TbRhGe, collinear with the propagation vectors k = (\frac{1}{2}, 0, \frac{1}{2}) for HoRhGe and k = (0, \frac{1}{2}, 0) for ErRhGe. For ErRhGe, with an increasing temperature a change to the incommesurate sine modulated structure below T_{t} = 5 K is observed. For these compounds the Néel temperatures not obeyed the de Gennes function.

[figure: 13.4 cm]

Subject category :

4. Rare Earths and Actinides, Alloys and Compounds

Presentation mode : poster

Corresponding author : Prof. dr hab. Andrzej Szytula

Address for correspondence : Instytut Fizyki UJ 30-059 Krakow, Reymonta 4 tel (12) 632 48 88 / 5546 fax: (12) 633 70 86

Email address : szytula@if.uj.edu.pl